# Subspace State Transfer in Coined Quantum Walks

### v1

@@ 5 $\gamma$-strong-cospectrality
In this section, we provide a characterization for $\gamma$-strong-cospectrality using certain rational functions. Much of the theory can be seen as an extension of [10] from two vertices to two sets of vertices. We will use these results to develop a polynomial-time algorithm that tests whether perfect pointwise subspace state transfer occurs at an integer step, given the subspace and coins are rational, in Section 7.
@@
When $S=T$, we will abbreviate $\psi_{S,T}(x)$ as $\psi_{S}(x)$. Since $E_{\lambda}$ is positive semidefinite, $E_{\lambda}D_{S}=0$ if and only if $\operatorname{tr}((E_{\lambda})_{S,S})=0$. Thus the eigenvalue support of $S$ can be read off from the poles of $\psi_{S}(x)$.
@@
Let $Y$ be a Hermitian weighted digraph. Let $S$ and $T$ be two subsets of vertices in $Y$ of equal size. Then $S$ and $T$ are $\gamma$-strongly-cospectral for some unimodular complex number $\gamma$ if and only if the following hold.
@@
Moreover, if $S$ and $T$ are $\gamma$-strongly-cospectral, then $\Lambda_{S,T}^{\gamma}$ is given by Equation (9), and $\Lambda_{S,T}^{-\gamma}$ is given by Equation (10).
@@
Suppose first that $S$ and $T$ are $\gamma$-strongly-cospectral. Then $\left\lvert\gamma\right\rvert=1$, and for each $\lambda$,
$E_{\lambda}D_{S}=\pm\gamma E_{\lambda}D_{T}.$ (11)
@@
$\operatorname{tr}((E_{\lambda})_{S,S})=\langle E_{\lambda}D_{S},E_{\lambda}D_{S}\rangle=\langle E_{\lambda}D_{T},E_{\lambda}D_{T}\rangle=\operatorname{tr}((E_{\lambda})_{T,T})$
for each $\lambda$. This shows (i). Now multiply both sides of Equation (11) on the left by $D_{S}$ and take the trace. We obtain
$\operatorname{tr}((E_{\lambda})_{S,S})=\operatorname{tr}(D_{S}E_{\lambda}D_{S})=\pm\gamma\operatorname{tr}(D_{S}E_{\lambda}D_{T})=\pm\gamma\operatorname{tr}((E_{\lambda})_{S,T}).$ (12)
@@
$\left\lvert\langle E_{\lambda}D_{S},E_{\lambda}D_{T}\rangle\right\rvert^{2}\leq\langle E_{\lambda}D_{S},E_{\lambda}D_{S}\rangle\langle E_{\lambda}D_{T},E_{\lambda}D_{T}\rangle,$
with equality if and only if $E_{\lambda}D_{S}$ is a scalar multiple of $E_{\lambda}D_{T}$, it follows that
$E_{\lambda}D_{S}=\pm\gamma E_{\lambda}D_{T}.$
Therefore, $S$ and $T$ are $\gamma$-strongly-cospectral.
@@
Let $Y$ be a Hermitian weighted digraph with Hermitian adjacency matrix $H$. Let $S$ and $T$ be two subsets of vertices in $Y$ with equal size. Suppose there is a diagonal matrix $\Delta$ such that both $\Delta H\Delta^{-1}$ and $\Delta_{S,S}\Delta^{-1}_{T,T}$ have rational entries. If $S$ and $T$ are $\gamma$-strongly-cospectral, then they are strongly-cospectral, and $\Lambda_{S,T}^{+}$ and $\Lambda_{S,T}^{-}$ are each closed under algebraic conjugates.
@@ 6 Rational subspaces
which is diagonally similar to a matrix with rational entries. If $S$ and $T$ are the $W$-clones of $a$ and the $W$-clones of $b$, then
@@
Proof. Setting $a=b$ and $W=V$ in Theorem 4.1 shows that $U$ is pointwise $W$-periodic with integer period $\tau$ if and only if for each $\lambda\in\Lambda_{S}$,
@@
7 Pointwise subspace state transfer
Sometimes we may want to preserve the amplitude on each arc during state transfer. We say there is perfect pointwise subspace state transfer from $a$ to $b$ if there is some time $t\in{\mathbb{R}}$ and some subspace $W\in\operatorname{col}(C_{a}+I)\cap\operatorname{col}(C_{b}+I)$ such that for any unit vector $w\in W$, there is a unimodular $\gamma$ such that
$U^{t}x_{a}(w)=\gamma x_{b}(w).$
When $W$ is known, this is called perfect pointwise $W$-transfer from $a$ to $b$. It follows from linearity of $U$ that if this happens, then the phase $\gamma$ does not depend on the choice of $w$.
We are primarily interested in pointwise $W$-transfer at integer steps where both $C$ and $W$ are rational. The next result shows this can be tested in polynomial time.
@@ 7 Pointwise subspace state transfer
Let $C$ be a rational coin and $W$ a rational subspace of $\operatorname{col}(C_{a}+I)\cap\operatorname{col}(C_{b}+I)$. The following algorithm tests whether pointwise $W$-transfer from $a$ to $b$ occurs at integer steps and, if this happens, returns the minimum step needed for such transfer. Moreover, the running time of this algorithm is polynomial in the number of vertices of the original graph.
@@
    Check if $\psi_{S}(x)=\psi_{T}(x)$. If not, pointwise $W$-transfer does not occur from $a$ to $b$.
@@
    Check if $g^{\sharp}(x)$ factors into cyclotomic polynomials. If not, pointwise $W$-transfer does not occur from $a$ to $b$ at any integer step. Otherwise, let $L$ be the set of positive integers such that
@@
    Check if $\tau$ is even. If not, pointwise $W$-transfer does not occur from $a$ to $b$ at any integer step. Otherwise, let
@@
    $g^{\sharp}_{+}(x)=\prod_{m\in L_{+}}\Phi_{m}(x),\quad g^{\sharp}_{-}(x)=\prod_{m\in L_{-}}\Phi_{m}(x).$
    If so, pointwise $W$-transfer occurs from $a$ to $b$ at time $\tau/2$. Otherwise, it does not occur at any integer step.
Proof. By Theorem 4.1 and Theorem 6.2, it suffices to show that Steps (5) to (6) test strong cospectrality of $S,T$ and the parity condition for $\Lambda_{S,T}^{\pm}$ in polynomial time. Indeed, as $C$ and $W$ are both rational, $S$ and $T$ are strongly cospectral if and only if each irreducible factor of $g^{\sharp}(x)$ divides exactly one of $g^{\sharp}_{+}(x)$ and $g^{\sharp}_{-}(x)$. Given $U$ is $W$-periodic at $a$ with mininum integer period $\tau$, pointwise $W$-transfer occurs from $a$ to $b$ at time $\tau/2$ if and only if $\tau$ is even, and for each $m\in L$, the cyclotomic polynomial $\Phi_{m}(x)$ divides $g^{\sharp}_{+}(x)$ if and only if $\tau/m$ is even.
@@ 8 Transfer between marked vertices
Since most coins in this walk are reflections about one-dimensional subspaces, we can afford a a characterization of perfect pointwise subspace state transfer between $a$ and $b$ that involves only the coins $C_{a}$ and $C_{b}$ rather than their specific decompositions into $K$ and $L$. To start, let $M$ be the block diagonal matrix
@@
Note that $\operatorname{col}(M)$ is $G$-invariant:
$GM=MH,$
and so the eigenvectors for $H$ and $G$ are related in the following way.
8.1 Lemma.
Let $H$ be the Hermitian weighted digraph relative to any orthonormal basis induced by $W$. Let $M$ be the matrix with this basis as columns. Let $G=MHM^{*}$.
  1. (i)
    If $y$ is a $\lambda$-eigenvector for $H$, then $My$ is a $\lambda$-eigenvector for $G$.
  2. (ii)
    If $z$ is a $\lambda$-eigenvector for $G$, then either $M^{*}z=0$, in which case $\lambda=0$, or $M^{*}z$ is a $\lambda$-eigenvector for $H$.
In other words, if
@@
is the spectral decomposition of $H$. As a result, for any subspace $W$ of $\operatorname{col}(C_{a}+I)$ and the set $S$ of $W$-clones relative to $H$, the eigenvalue support $\Lambda_{S}$ consists of eigenvalues $\lambda$ for which the projection of the column space of $E_{\lambda}[\operatorname{cl}(a),\operatorname{cl}(a)]$ onto $W$ is nonzero. Since $G$ is independent of the basis for $W$, we will also refer to this set of eigenvalues as the eigenvalue support of $x_{a}(W)$, and denote it by $\Lambda_{a(W)}$:
$\Lambda_{a(W)}=\{\lambda:\operatorname{proj}_{W}\!\left(E_{\lambda}[\operatorname{cl}(a),\operatorname{cl}(a)]\right)\neq 0\}.$
Given two subspaces $W\subseteq\operatorname{col}(C_{a}+I)$ and $V\subseteq\operatorname{col}(C_{b}+I)$, we say $x_{a}(W)$ and $x_{b}(V)$ are $\gamma$-strongly-cospectral for some unimodular $\gamma$ if for each eigenprojection $E_{\lambda}$ of $G$, we have
$\operatorname{proj}_{W}\!\left(E_{\lambda}[\operatorname{cl}(a),\operatorname{cl}(a)]\right)=\pm\gamma\operatorname{proj}_{V}\!\left(E_{\lambda}[\operatorname{cl}(b),\operatorname{cl}(b)]\right),$
and we define
$\Lambda_{a(W),b(V)}^{\gamma}=\{\lambda:\operatorname{proj}_{W}\!\left(E_{\lambda}[\operatorname{cl}(a),\operatorname{cl}(a)]\right)=\gamma\operatorname{proj}_{V}\!\left(E_{\lambda}[\operatorname{cl}(b),\operatorname{cl}(b)]\right)\}$
@@
$\Lambda_{a(W),b(V)}^{-\gamma}=\{\lambda:\operatorname{proj}_{W}\!\left(E_{\lambda}[\operatorname{cl}(a),\operatorname{cl}(a)]\right)=-\gamma\operatorname{proj}_{V}\!\left(E_{\lambda}[\operatorname{cl}(b),\operatorname{cl}(b)]\right).$
This, together with Theorem 4.1, leads to a basis-free characterization of perfect subspace state transfer between marked vertices.
@@
Let $U=RC$ be the transition matrix of a Grover coined walk on $X$ with marked vertices $a$ and $b$. Let $W\subseteq\operatorname{col}(C_{a}+I)$ and $V\subseteq\operatorname{col}(C_{b}+I)$ be two subspaces. Let
@@
be the spectral decomposition of the $(C_{a},C_{b})$-blow-up of $X$. There is perfect $(W,V)$-transfer from $a$ to $b$ at time $t$ if and only if the following hold.
@@
    $x_{a}(W)$ and $x_{b}(V)$ are $\gamma$-strongly-cospectral.
@@
    If $\lambda\in\Lambda_{a(W),b(V)}^{\gamma}$, then
    $\lambda=\cos(j\pi/t)$
    for some even integer $j$.
@@
    If $\lambda\in\Lambda_{a(W),b(V)}^{\gamma}$, then
    $\lambda=\cos(j\pi/t)$
    for some odd integer $j$.

### v2

@@ 6 $\gamma$-strong-cospectrality
In this section, we provide a characterization for $\gamma$-strong-cospectrality using certain rational functions. Much of the theory can be seen as an extension of [12] from two vertices to two sets of vertices. We will use these results to develop a polynomial-time algorithm that tests whether pointwise perfect subspace state transfer occurs at an integer step, given the subspace and coins are rational, in Section 7. This complements the polynomial-time algorithm for perfect state transfer in continuous quantum walks [11].
@@
When $S=T$, we will abbreviate $\psi_{S,T}(x)$ as $\psi_{S}(x)$. Since $E_{\lambda}$ is positive semidefinite, $E_{\lambda}B_{S}=0$ if and only if $\operatorname{tr}((E_{\lambda})_{S,S})=0$. Thus the eigenvalue support of $S$ can be read off from the poles of $\psi_{S}(x)$.
@@
Let $Y$ be a Hermitian weighted digraph. Let $S$ and $T$ be two subsets of vertices in $Y$ of equal size. Then $S$ is $\gamma$-strongly-cospectral to $T$ for some unimodular complex number $\gamma$ if and only if the following hold.
@@
Moreover, if $S$ is $\gamma$-strongly-cospectral to $T$, then $\Lambda_{S,T}^{\gamma}$ is given by Equation (11), and $\Lambda_{S,T}^{-\gamma}$ is given by Equation (12).
@@
Suppose first that $S$ is $\gamma$-strongly-cospectral to $T$. Then $\left\lvert\gamma\right\rvert=1$, and for each $\lambda$,
$E_{\lambda}B_{S}=\pm\gamma E_{\lambda}B_{T}.$ (13)
@@
$\operatorname{tr}((E_{\lambda})_{S,S})=\langle E_{\lambda}B_{S},E_{\lambda}B_{S}\rangle=\langle E_{\lambda}B_{T},E_{\lambda}B_{T}\rangle=\operatorname{tr}((E_{\lambda})_{T,T})$
for each $\lambda$. This shows (i). Now multiply both sides of Equation (13) on the left by $B_{S}$ and take the trace. We obtain
$\operatorname{tr}((E_{\lambda})_{S,S})=\operatorname{tr}(B_{S}E_{\lambda}B_{S})=\pm\gamma\operatorname{tr}(B_{S}E_{\lambda}B_{T})=\pm\gamma\operatorname{tr}((E_{\lambda})_{S,T}).$ (14)
@@
$\left\lvert\langle E_{\lambda}B_{S},E_{\lambda}B_{T}\rangle\right\rvert^{2}\leq\langle E_{\lambda}B_{S},E_{\lambda}B_{S}\rangle\langle E_{\lambda}B_{T},E_{\lambda}B_{T}\rangle,$
with equality if and only if $E_{\lambda}B_{S}$ is a scalar multiple of $E_{\lambda}B_{T}$, it follows that
$E_{\lambda}B_{S}=\pm\gamma E_{\lambda}B_{T}.$
Therefore, $S$ is $\gamma$-strongly-cospectral to $T$.
@@
Let $Y$ be a Hermitian weighted digraph with Hermitian adjacency matrix $H$. Let $S$ and $T$ be two subsets of vertices in $Y$ with equal size. Suppose there is a diagonal matrix $\Delta$ such that both $\Delta H\Delta^{-1}$ and $\Delta_{S,S}\Delta^{-1}_{T,T}$ have rational entries. If $S$ is $\gamma$-strongly-cospectral to $T$, then they are strongly-cospectral, and $\Lambda_{S,T}^{+}$ and $\Lambda_{S,T}^{-}$ are each closed under algebraic conjugates.
@@ 7 Rational subspaces
which is diagonally similar to a matrix with rational entries. If $S$ and $T$ are the $W$-clones of $a$ and the $V$-clones of $b$, then
@@
Proof. By Theorem 4.3, $U$ is pointwise $W$-periodic with integer period $\tau$ if and only if for each $\lambda\in\Lambda_{S}$,
@@
We are primarily interested in pointwise perfect $W$-transfer at integer steps where both $C$ and $W$ are rational. Using Theorem 7.2 and results from Section 6,, we show this can be tested in polynomial time.
@@
Let $C$ be a rational coin and $W$ a rational subspace of $\operatorname{col}(C_{a}+I)\cap\operatorname{col}(C_{b}+I)$. The following algorithm tests whether pointwise perfect $W$-transfer from $a$ to $b$ occurs at integer steps and, if this happens, returns the minimum step needed for such transfer. Moreover, the running time of this algorithm is polynomial in the number of vertices of the original graph.
@@
    Check if $\psi_{S}(x)=\psi_{T}(x)$. If not, pointwise perfect $W$-transfer does not occur from $a$ to $b$.
@@
    Check if $g^{\sharp}(x)$ factors into cyclotomic polynomials. If not, pointwise perfect $W$-transfer does not occur from $a$ to $b$ at any integer step. Otherwise, let $L$ be the set of positive integers such that
@@
    Check if $\tau$ is even. If not, pointwise perfect $W$-transfer does not occur from $a$ to $b$ at any integer step. Otherwise, let
@@
    $g^{\sharp}_{+}(x)=\prod_{m\in L_{+}}\Phi_{m}(x),\quad g^{\sharp}_{-}(x)=\prod_{m\in L_{-}}\Phi_{m}(x),$
    or
    $g^{\sharp}_{+}(x)=\prod_{m\in L_{-}}\Phi_{m}(x),\quad g^{\sharp}_{-}(x)=\prod_{m\in L_{+}}\Phi_{m}(x).$
    If so, pointwise perfect $W$-transfer occurs from $a$ to $b$ at time $\tau/2$. Otherwise, it does not occur at any integer step.
Proof. By Theorem 7.2, the quantum walk passes Step (3) and (4) if and only if $U$ is $W$-periodic at $a$ with mininum integer period $\tau$. Thus the minimum time for pointwise perfect $W$-transfer to occur from $a$ to $b$ is $\tau/2$, which is an integer if and only if $\tau$ is even.
Now suppose $U$ is pointwise $W$-periodic at $a$ with even minimum period $\tau$. Let $t=\tau/2$. By Theorem 5.2 and Corollary 6.4, it suffices to show that Steps (2), (5) and (6) test strong cospectrality between $S$ and $T$ and the parity condition for $\Lambda_{S,T}^{\pm}$ in polynomial time. Indeed, as $C$ and $W$ are both rational, $S$ and $T$ are strongly cospectral if and only if they are cospectral and each irreducible factor of $g^{\sharp}(x)$ divides exactly one of $g^{\sharp}_{+}(x)$ and $g^{\sharp}_{-}(x)$. Note that if $\tau/m$ is even, then all primitive $m$-th roots unity are also $t$-th roots of unity, and if $\tau/m$ is odd, then no primitive $m$-th root of unity is a $t$-th root of unity. Therefore, we have pointwise perfect $W$-transfer in the form
$U^{t}x_{a}(w)=x_{b}(w),\quad\forall w\in W$
if and only if
$g^{\sharp}_{+}(x)=\prod_{m\in L_{+}}\Phi_{m}(x),\quad g^{\sharp}_{-}(x)=\prod_{m\in L_{-}}\Phi_{m}(x),$
and pointwise perfect $W$-transfer in the form
$U^{t}x_{a}(w)=-x_{b}(w),\quad\forall w\in W$
if and only if
$g^{\sharp}_{+}(x)=\prod_{m\in L_{-}}\Phi_{m}(x),\quad g^{\sharp}_{-}(x)=\prod_{m\in L_{+}}\Phi_{m}(x).$
@@ 8 Transfer between marked vertices
Since most coins in this walk are reflections about one-dimensional subspaces, we can afford a characterization of pointwise perfect subspace state transfer between $a$ and $b$ that involves only the coins $C_{a}$ and $C_{b}$ rather than their specific decompositions into $K$ and $L$. To start, let $M$ be the block diagonal matrix
@@
Note that $\operatorname{col}(M)$ is $G$-invariant. Thus if
@@
is the spectral decomposition of $H$. As a result, for any subspace $W$ of $\operatorname{col}(C_{a}+I)$ and the set $S$ of $W$-clones relative to $H$, the eigenvalue $\lambda$ of $H$ lies in $\Lambda_{S}$ if and only if the column space of $E_{\lambda}[\operatorname{cl}(a),\operatorname{cl}(a)]$ is not orthogonal to $W$. Since $G$ is independent of the basis for $W$, we will also refer to $\Lambda_{S}$ as the eigenvalue support of $x_{a}(W)$, and denote it by $\Lambda_{a(W)}$:
$\Lambda_{a(W)}=\{\lambda:E_{\lambda}[\operatorname{cl}(a),\operatorname{cl}(a)]\upharpoonright_{W}\neq 0\}.$
Moreover, given $W\subseteq\operatorname{col}(C_{a}+I)\cap\operatorname{col}(C_{b}+I)$, the $W$-clones of $a$ are $\gamma$-strongly-cospectral to the $W$-clones of $b$ if and only if
$E_{\lambda}[\operatorname{cl}(a),\operatorname{cl}(a)\cup\operatorname{cl}(b)]\upharpoonright_{W}=\pm\gamma E_{\lambda}[\operatorname{cl}(b),\operatorname{cl}(a)\cup\operatorname{cl}(b)]\upharpoonright_{W}$
for each eigenvalue $\lambda$ of $G$. When this happens, we say $x_{a}(W)$ is $\gamma$-strongly cospectral to $x_{b}(V)$, and define
$\Lambda_{a(W),b(W)}^{\gamma}=\{\lambda:E_{\lambda}[\operatorname{cl}(a),\operatorname{cl}(a)\cup\operatorname{cl}(b)]\upharpoonright_{W}=\gamma E_{\lambda}[\operatorname{cl}(b),\operatorname{cl}(a)\cup\operatorname{cl}(b)]\upharpoonright_{W}\}$
@@
$\Lambda_{a(W),b(W)}^{-\gamma}=\{\lambda:E_{\lambda}[\operatorname{cl}(a),\operatorname{cl}(a)\cup\operatorname{cl}(b)]\upharpoonright_{W}=-\gamma E_{\lambda}[\operatorname{cl}(b),\operatorname{cl}(a)\cup\operatorname{cl}(b)]\upharpoonright_{W}\}.$
This, together with Theorem 5.2 and Theorem 5.3, leads to basis-free characterizatiosn of pointwise perfect and pretty good $W$-transfer between marked vertices.
@@
Let $U=RC$ be the transition matrix of a Grover coined walk on $X$ with marked vertices $a$ and $b$. Let $W\subseteq\operatorname{col}(C_{a}+I)\cap\operatorname{col}(C_{b}+I)$ be a subspace. Let
@@
be the spectral decomposition of the $(C_{a},C_{b})$-blow-up of $X$. There is pointwise perfect $W$-transfer from $a$ to $b$ at time $t$ if and only if the following hold for some unimodular $\gamma$.
@@
    $x_{a}(W)$ is $\gamma$-strongly-cospectral to $x_{b}(W)$.
@@
    $t\arccos\lambda$ is an even multiple of $\pi$ for each eigenvalue $\lambda\in\Lambda_{a(W),b(W)}^{\gamma}$.
@@
    $t\arccos\lambda$ is an odd multiple of $\pi$ for each eigenvalue $\lambda\in\Lambda_{a(W),b(W)}^{-\gamma}$.
8.2 Theorem.
Let $U=RC$ be the transition matrix of a Grover coined walk on $X$ with marked vertices $a$ and $b$. Let $W\subseteq\operatorname{col}(C_{a}+I)\cap\operatorname{col}(C_{b}+I)$ be a subspace. Let
$G=\sum_{\lambda}\lambda E_{\lambda}$
be the spectral decomposition of the $(C_{a},C_{b})$-blow-up of $X$. There is pointwise pretty good $W$-transfer from $a$ to $b$ if and only if the following hold for some unimodular $\gamma$.
  1. (i)
    $x_{a}(W)$ is $\gamma$-strongly-cospectral to $x_{b}(W)$.
  2. (ii)
    For any set $\{\ell_{\lambda}:\lambda\in\Lambda_{a(W)}\}$ of integers such that
    $\sum_{\lambda\in\Lambda_{a(W)}}\ell_{\lambda}\arccos\lambda=0,$ (15)
    we have
    $\sum_{\lambda\in\Lambda^{-\gamma}_{a(W),b(W)}}\ell_{\lambda}\equiv 0\pmod{2}$.
Moreover, the target states can be approximated in integer steps if and only if (i) and (ii) hold with Equation (15) replaced by
$\sum_{\lambda\in\Lambda_{a(W)}}\ell_{\lambda}\arccos\lambda\equiv 0\pmod{2\pi}$.
When an explicit eigenbasis for $G$ is known, we may work with the following equivalent definitions of eigenvalue support and $\gamma$-strong-cospectrality.
8.3 Lemma.
Let $U=RC$ be the transition matrix of a Grover coined walk on $X$ with marked vertices $a$ and $b$. Let $W\subseteq\operatorname{col}(C_{a}+I)\cap\operatorname{col}(C_{b}+I)$ be a subspace.
  1. (i)
    An eigenvalue $\lambda$ of $G$ lies in $\Lambda_{a(W)}$ if and only if some $\lambda$-eigenvector $z$ satisfies $\operatorname{proj}_{W}\!\left(z[\operatorname{cl}(a)]\right)\neq 0$.
  2. (ii)
    $x_{a}(W)$ is $\gamma$-strongly-cospectral to $x_{b}(W)$ if and only if for each eigenprojection $E_{\lambda}$ of $G$, there is $\sigma_{\lambda}\in\{1,-1\}$ such that every $\lambda$-eigenvector $z$ satisfies
    $\operatorname{proj}_{W}\!\left(z[\operatorname{cl}(a)]\right)=\sigma_{\lambda}\gamma\operatorname{proj}_{W}\!\left(z[\operatorname{cl}(b)]\right).$
Proof. The results follow from the fact that $E_{\lambda}$ is a sum of outer products of any orthonormal basis for the $\lambda$-eigenspace.
@@
Proof. Expanding $Gz=\lambda z$ gives
$\begin{cases}F^{*}y=\lambda x,\\ Fx+By=\lambda y.\end{cases}$ (16)
If $\lambda=0$, the above is equivalent to
$F^{*}y=0,\quad Fx+By=0.$
Otherwise, multiplying the second equation by $\lambda$ and plugging the first equation in (16) yields
$(\lambda^{2}I-\lambda B-FF^{*})y=0.$
Conversely, if
$\lambda\neq 0,\quad F^{*}y=\lambda x,\quad(\lambda^{2}I-\lambda B-FF^{*})y=0,$
then it is not hard to see that both equations in (16) hold.